OpenStudy (anonymous):

Fan & Medals!! 1. Which rule describes the translation left 4 units and up 5 units? 1. Which rule describes the translation left 4 units and up 5 units? A.(x, y) (x – 5, y + 4) B.(x, y) (x + 4, y - 5) C.(x, y) (x – 4, y + 5) D.(x, y) (x + 5, y – 4)

OpenStudy (anonymous):

i would say C

OpenStudy (anonymous):

x is tending toward left so as the rule of translation x will x+a = x+4 y is tending toward upward so as the rule of translation y will y+a = y+5 http://www.purplemath.com/modules/fcntrans2.htm